Solve the following system of equations analytically. y = (x + 4; y = 8 – x
Answer #1

Given y = 5x +4 y = 8-2 By equating above two equations Jx+4=8-2 By squaring on both sides =;} = (8-x) (5x+4 X+4 = 64 tx – 1